OpenStudy (anonymous):

Omar paved a 6 foot walkway using 20 stones. He now wants to extend the walkway to 9 feet. How many more stones will he need to buy?

OpenStudy (shiraz14):

Number of stones needed = [(9-6)/6] x 20 = (20/2) stones = 10 stones

OpenStudy (yanasidlinskiy):

Set up a proportion: 6/20 = 9/x x will be the total number of stones. Then perform x-20 to get the number of additional stones to buy. Notice that x is the number of stones needed for 9 ft, Since we already had stones for 6 ft, => We only need x - 20 stones MORE :)